OpenStudy (anonymous):

1 + sec^2(x) sin^2(x) = sec^2(x)

OpenStudy (anonymous):

please show steps

OpenStudy (anonymous):

\[\begin{align*}1+\sec^2x\sin^2x&=\sec^2x\left(\frac{1}{\sec^2x}+\sin^2x\right)\\\\ &=\sec^2x\left(\cos^2x+\sin^2x\right)\\ &=\sec^2x(1)\\ &=\sec^2x\end{align*}\]

OpenStudy (anonymous):

how did you get 17sec^2(X)

OpenStudy (anonymous):

1/sec^2(x)

OpenStudy (anonymous):

Anything divided by itself is 1, right? \[1=\frac{\sec^2x}{\sec^2x}=\sec^2x\left(\frac{1}{\sec^2x}\right)\]

OpenStudy (anonymous):

ok i guess

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

Is 2 times 1/2 equal to 1? The answer's yes. What I did in the first step was work backwards - factorizing. \[1=\frac{2}{2}=2\left(\frac{1}{2}\right)\] Similarly, \[1=\frac{\sec^2x}{\sec^2x}=\sec^2x\left(\frac{1}{\sec^2x}\right)\] So, the \(1\) and \(\sec^2x\sin^2x\) can be written so that they share a common factor: \[\begin{align*}1+\color{blue}{\sec^2x}\sin^2x&=\frac{\color{blue}{\sec^2x}}{\sec^2x}+\color{blue}{\sec^2x}\sin^2x\\ &=\color{blue}{\sec^2x}\left(\frac{1}{\sec^2x}+\sin^2x\right)\end{align*}\]

OpenStudy (anonymous):

OMG THANK YOU SO MUCH

OpenStudy (anonymous):

wait how did you get 1/sec^2(x) + sin^2(x) where did that addition sign come from

OpenStudy (anonymous):

The plus is on the left side of the equation: \[1\color{red}+\sec^2x\sin^2x\]